OpenStudy (anonymous):

Plz check my answer ``Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 8x + 11y = -50 -32x – 44y = -200

OpenStudy (anonymous):

8x + 11y = -50 -32x – 44y = -200 ----- Multiply thru the 1st equation by -4 --- -32x - 44y = 200 -32x - 44y = -200 --------------- Subtract: 0 = 400 --- That is a contradiction. There is no ordered pair (x,y) that satisfies the system. The lines are parallel.

OpenStudy (anonymous):

looks right

OpenStudy (anonymous):

ty